A certain class of lamp takes 0.45A at 220V. How many such lamps can be run for an expenditure of power equal to 1 1/3 horsepower?

1
Expert's answer
2022-02-02T14:39:37-0500

Power

P=IV

Now power

"P=0.45\times220=99W"

"P'=\frac{11}{3}" Hp

"P'=\frac{11\times746}{3}=2735.33W"

Number of lamp="\frac{P'}{P}"

"n=\frac{2735.33}{99}=27.62"

"n=27" (Approximately)